Heres the problem, Sergio sells a mixture or Virginia peanuts and Spanish peanuts for $3.40 per pound. To make the mixture, he uses Virginia peanuts that cost $3.50 per pound and Spanish peanuts that cost $3 per pund. He mixes 10 pounds at a time. (a)How many pounds or Virginia peanuts does Sergio use? (b)How many pounds Spanish peanuts does Sergio use?
Answer by ankor@dixie-net.com(22740)

Here's the problem, Sergio sells a mixture of Virginia peanuts and Spanish peanuts for $3.40 per pound. To make the mixture, he uses Virginia peanuts that cost $3.50 per pound, and Spanish peanuts that cost $3 per pound. He mixes 10 pounds at a time.
:
(a)How many pounds of Virginia peanuts does Sergio use?
Assume he is making 10 lb of the $3.40 mixture
:
Let x = no. of pounds of the Virginia ($3.50) peanuts
Then (10-x) = no. of pounds of the Spanish ($3) peanuts
:
The Equation:
3.5x + 3(10-x) = 3.4(10)
3.5x + 30 - 3x = 34
.5x = 34 - 30
x = 4/.5
x = 8 pounds of Virginia peanuts
:
(b)How many pounds Spanish peanuts does Sergio use?
Spanish peanuts must be 2 pounds since the total is 10 pounds
:
Check:
3.5(8) + 3(2) = 3.4(10)
28 + 6 = 34